OpenStudy (anonymous):

how do i find the distance traveled in 15 seconds by an object moving with a velocity of 20 feet per second

OpenStudy (***[isuru]***):

It's really simple U have to use a linear motion equation here hence there are no acceleration S = ut ( S= displacement/distance , U = velocity , t = time) S = 20 * 15 s = 300 feet So the answer is 300 feet which is 91.44 meter's. Hope this will help u!

OpenStudy (anonymous):

It does, thanks you!

OpenStudy (***[isuru]***):

u r welcome!

OpenStudy (anonymous):

what about when the object moves with a velocity of v(t)=20+7cos(t) ft/sec?

OpenStudy (anonymous):

do I plug in 15 into the function for t?

OpenStudy (***[isuru]***):

If the velocity is given by the function v(t)=20+7cos(t) [ where t = time] u have to substitute the time given in the question , inside to the function v(t) and obtain a value for the velocity. Then u can use the value u got for velocity in any linear motion equation